Shift report 20030204 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Feb 04 2003 - 07:52:36 EST

  • Next message: Christian Holm Christensen: "My Master Thesis"
    ------------------------------------------------------------
    Date: Tue Feb  4 07:52:35 2003, Shift: 00:00-08:00
    Supervisor: IA, with: SZ
    Log book#13 pages 146 - 147
    
    Shift Summary:
    ------------------------------------------------------------
    1. Several unsuccesfuly attempts to inject beam. At 5:45 we had good beam and start take data.
    2. Data runs:
       7774,7775 - problems with TPC's and DC's; bad runs.
       7776 - T2 off because of high background
       7777 - same T2 off
       7778 - T2 on
       Pedestal runs:
       7771 - DVM run
       7772 - TPC pedestal run
       7773 - Non-TPC pedestal run
    3. Problems with T2 probably because of high background oscilations from the begining of the store. Fleming called and solved the problem. T2 is now back on.
      Need more 214k trigger 6 events on FS 8 degrees 1/8B.
      We changed polarity in D5.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Tue Feb 04 2003 - 07:53:15 EST